OpenStudy (thats_kyy):

A system of equations is given below: y = –2x + 1 6x + 2y = 22 Which of the following steps could be used to solve by substitution?

OpenStudy (thats_kyy):

- 6x + 2 (-2x+1) = 22 - -2x + 1 = 6x + 2y - 6 (-2x + 1) + 2y = 22 - 6(y =-2x + 1

OpenStudy (thats_kyy):

@campbell_st @skittles_for_life6422

OpenStudy (thats_kyy):

@TheSmartOne Help Please

OpenStudy (thats_kyy):

@freckles

OpenStudy (thats_kyy):

someone please help me

OpenStudy (freckles):

I'm here.

OpenStudy (freckles):

One sec. Checking.

OpenStudy (thats_kyy):

okay

OpenStudy (freckles):

oh those are choices not your work :p

OpenStudy (freckles):

which choice do you think it is?

OpenStudy (thats_kyy):

a

OpenStudy (freckles):

a is awesome that method you choice is call solving the system by substitution y=-2x+1 so you enter this y into the next equation and yes you do get 6x + 2y = 22 that this becomes 6x+2(-2x+1)=22 :)

OpenStudy (thats_kyy):

so its right?

OpenStudy (freckles):

yes it is called solving the system by substitution as I said above :P